OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

baby girl hey merry xmast :))

OpenStudy (anonymous):

use the slope formula first.

OpenStudy (anonymous):

find the slope first so you do (2+4)/(-2) then you use that to plug it in to the formula y-y1=m(x-x1)

OpenStudy (anonymous):

yeah there yu go

OpenStudy (anonymous):

but i got -2/-2hat do i do then???

OpenStudy (anonymous):

for the slope? you should have gotten -3 then plug it in to y+2-=-3(x+2) then it would be y+2=-3x-6 then subtract the 2 on both sides so it would be y=-3x-8

OpenStudy (anonymous):

oh ok thank u baby girl :))

OpenStudy (anonymous):

welcomeee haha